OpenStudy (anonymous):

The price of grapes at a farm is $1.85 per pound. Which equation can be used to determine c, the total price of n pounds of grapes?

OpenStudy (rainbow_rocks03):

Is there a picture?

OpenStudy (rainbow_rocks03):

or answers?

OpenStudy (anonymous):

ill give you the options or equations to choose

OpenStudy (anonymous):

c= 1.85/n c = 1.85n c = 1.85 + n c = 1.85 - n

OpenStudy (rainbow_rocks03):

What do you think the answer is?

OpenStudy (anonymous):

c = 1.85 + n

OpenStudy (rainbow_rocks03):

Why do you think that is the answer?

OpenStudy (anonymous):

becuase its the total amount of money plus the total pounds of grapes

OpenStudy (rainbow_rocks03):

your answer is correct

OpenStudy (rainbow_rocks03):

i am 95% sure

OpenStudy (e.mccormick):

When you have some quantity at a particular value, you multiply to get the total. c = 1.85 + n is the price of one poinf plus the number of pounds. That is not multiplication